# Spectral Text Fusion: A Frequency-Aware Approach to Multimodal Time-Series Forecasting

### v1

@@ 1 Introduction
Although recent efforts have been made to incorporate multi-modality into time series forecasting, existing approaches remain limited. MM-TSFlib (Liu et al., 2024a) employs a late-fusion architecture that views the whole text sequence as a vector by a separate pipeline with time series before combining outputs with a simple linear weighting mechanism, neglecting the cross-modal interaction. Meanwhile, TaTS (Li et al., 2025) relies on the text-periodicity assumption to align text with time series by treating time series-paired texts as auxiliary variables of the time series itself. We argue that they still cannot exploit the textual information well, which frequently encodes contextual signals that simultaneously influence multiple temporal scales. For instance, a central bank policy announcement may trigger immediate market volatility (high-frequency components) while also initiating a prolonged economic trend (low-frequency components). Similarly, news of supply chain disruptions might cause both short-term price spikes and longer-term inventory adjustments.
@@ 2 Related Works
Multimodal time series forecasting with textual data. Multimodal approaches address the limitations of unimodal methods by integrating complementary data sources (Jiang et al., 2025). Text2Freq (Lo et al., 2024) pioneers frequency-domain alignment but it utilizes textual data that essentially serves as a description or representation of the time series itself; the text is closely tied to the numerical data and does not provide additional or external contextual information. MM-TSFlib (Liu et al., 2024a) linearly combines the predictions from text and time series from separate models, neglecting the cross-modal interaction. TaTS (Li et al., 2025) projects text embeddings as auxiliary time series variables but conflates their influence across frequencies. TimeXer (Wang et al., 2024b) represents a significant advancement by empowering canonical Transformers, employing a global token mechanism to bridge information between exogenous and endogenous variables by using the attention mechanism at the temporal domain. While effective for general information transfer, this approach treats the influence of exogenous variables uniformly across all temporal dynamics. Compared to TaTS or TimeXer, which use time-domain fusion, our SpecTF, using cross-attention in the frequency domain, explicitly models the influence of textual data at multiple temporal scales through different frequency components.
@@ 3.1 Multimodal Time Series Forecasting
In the context of multimodal time series forecasting, the primary goal is to predict future values of a time series while utilizing additional modalities that are temporally paired with the time series. This paper focuses specifically on text-enhanced forecasting. In particular, given a collection of multivariate time series samples and textual data with a lookback window $L$, the number of variables $N$, we denote the input as pairs $(x_{1},s_{1}),\ldots,(x_{L},s_{L})$, where $X^{h}=\left[x_{1},\ldots,x_{L}\right]\in\mathbb{R}^{N\times L}$ is the numerical data and $S^{h}=\left[s_{1},\ldots,s_{L}\right]$ denotes the associated textual data. The objective of multimodal time series forecasting is to predict future numerical values $Y=\left[x_{L+1},\ldots,x_{L+H}\right]\in\mathbb{R}^{N\times H}$, where $H$ is the forecast horizon.
@@ 5 Experiments
Baselines. We compare our SpecTF with the representative and state-of-the-art models for both unimodal and multimodal time series forecasting. For unimodal forecasting, we compare SpecTF against FreTS (Yi et al., 2023), a frequency-based method. For multimodal time series forecasting, we include MM-TSF (Liu et al., 2024a) and TaTS (Li et al., 2025) with their best variant using iTransformer (Liu et al., 2024b) backbone for comparison. In addition, we also include TimeXer (Wang et al., 2024b) with text sequence as an exogenous variable by using our Text Embedding block rather than the exogenous embedding layer. For LLM-based models, we compare SpecTF against TimeLLM (jin2023time) and ChatTime (Wang_Qi_Wang_Sun_Zhuang_Wu_Zhang_Liao_2025).
@@ 5.1 Benchmarking Results
Table 1 presents the comparative results of SpecTF against all baselines across the nine domains from Time-MMD benchmark. For each dataset, we report the results averaged across three different seeds and all prediction lengths. Overall, SpecTF outperforms both unimodal and multimodal baselines in terms of MSE and MAE metrics on 8 out of 9 domains. When compared to the unimodal baseline FreTS and LLM-based approaches (TimeLLM, ChatTime), SpecTF demonstrates consistent improvements across all domains. This indicates that incorporating textual information in the frequency domain provides valuable additional context for forecasting. In comparison to the runner-up among multimodal approaches (TimeXer, MM-TSF, and TaTS), SpecTF still achieves superior performance in most domains, with the average improvement across 9 datasets of 3.82% and 2.25% in MSE and MAE, respectively. The most significant improvements over multimodal baselines are observed in the Energy, Environment, Health, and Security domains, where textual information often contains signals about regime shifts or policy changes that manifest at different temporal scales. However, SpecTF underperforms on the Economy dataset, where TaTS achieves marginally better results because the textual data contains more noise and shorter-term economic indicators that may be better suited to the time-domain alignment mechanism of TaTS.
@@ 5.2 Model Analysis
Model Efficiency. Table 4 shows that our SpecTF uses significantly fewer trainable parameters than TimeXer and far fewer than MM-TSF and TaTS. This leads to much lower computational costs, as shown by Multiply-Accumulate Operations (MACs) metric that counts the total number of multiplication and addition operations in a neural network. Besides, SpecTF achieves superior inference speed, outperforming all baselines. These results demonstrate that SpecTF improves efficiency without compromising performance, making it suitable for both resource-limited and large-scale applications.

### v2

@@ 1 Introduction
Although recent efforts have been made to incorporate multi-modality into time series forecasting, existing approaches remain limited. $M$-TSFlib (Liu et al., 2024a) employs a late-fusion architecture that views the whole text sequence as a vector by a separate pipeline with time series before combining outputs with a simple linear weighting mechanism, neglecting the cross-modal interaction. Meanwhile, TaTS (Li et al., 2025) relies on the text-periodicity assumption to align text with time series by treating time series-paired texts as auxiliary variables of the time series itself. We argue that they still cannot exploit the textual information well, which frequently encodes contextual signals that simultaneously influence multiple temporal scales. For instance, a central bank policy announcement may trigger immediate market volatility (high-frequency components) while also initiating a prolonged economic trend (low-frequency components). Similarly, news of supply chain disruptions might cause both short-term price spikes and longer-term inventory adjustments.
@@ 2 Related Works
Multimodal time series forecasting with textual data. Multimodal approaches address the limitations of unimodal methods by integrating complementary data sources (Jiang et al., 2025). Text2Freq (Lo et al., 2024) pioneers frequency-domain alignment but it utilizes textual data that essentially serves as a description or representation of the time series itself; the text is closely tied to the numerical data and does not provide additional or external contextual information. $M$-TSFlib (Liu et al., 2024a) linearly combines the predictions from text and time series from separate models, neglecting the cross-modal interaction. TaTS (Li et al., 2025) projects text embeddings as auxiliary time series variables but conflates their influence across frequencies. TimeXer (Wang et al., 2024b) represents a significant advancement by empowering canonical Transformers, employing a global token mechanism to bridge information between exogenous and endogenous variables by using the attention mechanism at the temporal domain. While effective for general information transfer, this approach treats the influence of exogenous variables uniformly across all temporal dynamics. Compared to TaTS or TimeXer, which use time-domain fusion, our SpecTF, using cross-attention in the frequency domain, explicitly models the influence of textual data at multiple temporal scales through different frequency components.
@@ 3.1 Multimodal Time Series Forecasting
In the context of multimodal time series forecasting, the primary goal is to predict future values of a time series while utilizing additional modalities that are temporally paired with the time series. This paper focuses specifically on text-enhanced forecasting. In particular, given a collection of multivariate time series samples and textual data with a lookback window $L$, the number of variables NN, we denote the input as pairs $(x_{1},s_{1}),\ldots,(x_{L},s_{L})$, where $X^{h}=\left[x_{1},\ldots,x_{L}\right]\in\mathbb{R}^{N\times L}$ is the numerical data and $S^{h}=\left[s_{1},\ldots,s_{L}\right]$ denotes the associated textual data. The objective of multimodal time series forecasting is to predict future numerical values $Y=\left[x_{L+1},\ldots,x_{L+H}\right]\in\mathbb{R}^{N\times H}$, where $H$ is the forecast horizon.
@@ 5 Experiments
Baselines. We compare our SpecTF with the representative and state-of-the-art models for both unimodal and multimodal time series forecasting. For unimodal forecasting, we compare SpecTF against FreTS (Yi et al., 2023), a frequency-based method. For multimodal time series forecasting, we include $M$-TSF (Liu et al., 2024a) and TaTS (Li et al., 2025) with their best variant using iTransformer (Liu et al., 2024b) backbone for comparison. In addition, we also include TimeXer (Wang et al., 2024b) with text sequence as an exogenous variable by using our Text Embedding block rather than the exogenous embedding layer. For LLM-based models, we compare SpecTF against TimeLLM (Jin et al., 2024) and ChatTime (Wang et al., 2025).
@@ 5.1 Benchmarking Results
Table 1 presents the comparative results of SpecTF against all baselines across the nine domains from Time-MMD benchmark. For each dataset, we report the results averaged across three different seeds and all prediction lengths. Overall, SpecTF outperforms both unimodal and multimodal baselines in terms of MSE and MAE metrics on 8 out of 9 domains. When compared to the unimodal baseline FreTS and LLM-based approaches (TimeLLM, ChatTime), SpecTF demonstrates consistent improvements across all domains. This indicates that incorporating textual information in the frequency domain provides valuable additional context for forecasting. In comparison to the runner-up among multimodal approaches (TimeXer, $M$-TSF, and TaTS), SpecTF still achieves superior performance in most domains, with the average improvement across 9 datasets of 3.82% and 2.25% in MSE and MAE, respectively. The most significant improvements over multimodal baselines are observed in the Energy, Environment, Health, and Security domains, where textual information often contains signals about regime shifts or policy changes that manifest at different temporal scales. However, SpecTF underperforms on the Economy dataset, where TaTS achieves marginally better results because the textual data contains more noise and shorter-term economic indicators that may be better suited to the time-domain alignment mechanism of TaTS.
@@ 5.2 Model Analysis
Model Efficiency. Table 4 shows that our SpecTF uses significantly fewer trainable parameters than TimeXer and far fewer than $M$-TSF and TaTS. This leads to much lower computational costs, as shown by Multiply-Accumulate Operations (MACs) metric that counts the total number of multiplication and addition operations in a neural network. Besides, SpecTF achieves superior inference speed, outperforming all baselines. These results demonstrate that SpecTF improves efficiency without compromising performance, making it suitable for both resource-limited and large-scale applications.